OpenStudy (anonymous):

what is the probability of drawing a blue card, not replacing it, and then drawing another blue card? (two red cards and three blue)

OpenStudy (debbieg):

So you have 2 red and 3 blue cards, you draw 2 without replacement, right? And you want the probably of both cards being blue.

OpenStudy (anonymous):

yes

OpenStudy (debbieg):

So, since these are dependent events (the outcome of drawing the first card changes the probability of the 2nd card being blue), you need to use: P(blue, blue) = P(1st card blue)*P(2nd blue|1st blue)

OpenStudy (debbieg):

So when you draw the first card, what is the probability that IT is blue?

OpenStudy (anonymous):

3/5

OpenStudy (debbieg):

You have 5 cards total, and 3 are blue. So what's the probability that you draw a blue card?

OpenStudy (debbieg):

Right! now, when you draw the 2nd card, what is the probabilitiy that IT is blue, GIVEN THAT the 1st card was blue?

OpenStudy (anonymous):

1/2?

OpenStudy (debbieg):

Right! Because there are 4 cards left, and 2 of them are blue. So now you need: P(blue, blue) = P(1st card blue)*P(2nd blue|1st blue) You have: P(1st card blue) = 3/5 P(2nd blue|1st blue) = 1/2 So what is P(blue, blue)?

OpenStudy (anonymous):

idk

OpenStudy (debbieg):

OK, what do you not understand? Did you read what I wrote above? Ask a question about whatever part you don't understand. That is how you will learn.

OpenStudy (debbieg):

do you know the probability formula for dependent events? When A and B are dependent, then: P(A and B)=P(A)*P(B|A)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the answer 3/10?

OpenStudy (debbieg):

Yes.

OpenStudy (anonymous):

thanks :)

OpenStudy (debbieg):

sure :)